OpenStudy (anonymous):

Calculate the average rate of change for the function f(x) = x^4 + 3x^3 - 5x^2 + 2x - 2, from x = -1 to x = 1. a. -1 b. 1 c. -5 d. 5

OpenStudy (anonymous):

\[\int\limits_{-1}^{1} (x^4 + 3x^3 -5x^2 + 2x -2 )/2\]

OpenStudy (anonymous):

I can't really tell what that means

OpenStudy (anonymous):

Have you taken integrals in calculus?

OpenStudy (anonymous):

No I'm in algebra 2

OpenStudy (anonymous):

\[ \frac{f(1)-f(-1)}{1-(-1)}=5 \]

OpenStudy (anonymous):

So the rate of change is 5.

OpenStudy (anonymous):

The rate of change of a function on an interval [a,b] is \[ \frac{f(b)-f(a)}{b-a} \]

OpenStudy (anonymous):

Yes it is 5